OpenStudy (anonymous):

A bag contains 5 white marbles, 4 black marbles, and 1 red marbles. Find the probability of picking a red marble?

OpenStudy (anonymous):

1 of 9?

OpenStudy (anonymous):

how many marbles are in the bag altogether

OpenStudy (anonymous):

10

OpenStudy (anonymous):

its 1 of 10

OpenStudy (anonymous):

do i have to simplify it to 1/5? or do i keep it as 1/10

OpenStudy (anonymous):

you cant simplify 1/10 any lower because of the 1

OpenStudy (anonymous):

A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick yellow marble?

OpenStudy (anonymous):

medal? lol

OpenStudy (anonymous):

i am getting 9/64 but that isn't one of the answers that comes up.. what am i doing wrong?

OpenStudy (anonymous):

1/8 9/56 3/28 1/28 these are the answers to chose from

OpenStudy (anonymous):

check and see if you have written your question correctly

OpenStudy (anonymous):

i did lol, but idk what I'm doing wrong

OpenStudy (anonymous):

it is simple addition, 5+4 is 9 and 9+1 is 10 their is 1 red marble so 1/10 marbles is red

OpenStudy (anonymous):

this is a different problem! haha

OpenStudy (anonymous):

A bag has 2 green marbles, 3 red marbles, and 3 yellow marbles. What is the probability that you pick a red marble, do not replace it, and pick yellow marble?

OpenStudy (anonymous):

i got the other one, but this one I'm getting 9/64 but that isn't an option choice i don't know what I'm doing wrong

OpenStudy (anonymous):

there are 8 marbles total 3 are red so that is 3/8 and u dont replace so now u got 7 3 are yellow so u have 3/7 now you multiply ill show u\[\frac{ 3 }{ 8 } \times \frac{ 3 }{ 7 }=\frac{ 9 }{ 56 }\]

OpenStudy (anonymous):

does 9/56 match one of your answers?

OpenStudy (anonymous):

ohh i see! i wasn't subtracting out the one you don't replace. i was doing 3/8 X 3/8 when it was supposed to be 3/8 X 3/7… Thats why i was getting 9/64! lol thank you:)

OpenStudy (anonymous):

and yes it does!

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

medal?

OpenStudy (anonymous):

Josie had a choice of 5 breads, 3 choices of meat, and 5 choices of cheeses to make a sandwich. How many different sandwiches can she make? when you add these all together you get 13.. but do i have to do anything else?

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

she has 75 different choices ill teach u a trick

OpenStudy (anonymous):

5(bread) times 3(meat) times 5(cheese) =75 different combos

OpenStudy (anonymous):

ohhhh i added, thats why. lol Thank you so much for you help! it helped me to be able to do my other problems

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

medal now? lolol

OpenStudy (anonymous):

this is my last one. What is the probability of rolling an even number on a six sided die and then rolling a 4? 2,4,6,8,10 are even numbers 3,5,7,9 are odd numbers a dice has 6 sides, so would it be 5/6 times something?

OpenStudy (anonymous):

lol sorry i didn't know what medal meant until just now haha